OpenStudy (anonymous):

find the exact value of cos(-7pi/4 radian)

OpenStudy (anonymous):

do you know what the \[\cos(\pi/4)=\]

OpenStudy (anonymous):

i know what the answe would be if wouldnt be negative but how would that change since it is negative

OpenStudy (anonymous):

you just have to find the specific quadrant it falls in 7/4=1.75 right so we go over one and a half radians in a negative direction 

OpenStudy (anonymous):

do you agree we end up in the first quadrant?

OpenStudy (anonymous):

so would the answer still be squareroot of 2/2 or would it change to either 1/2 squareroot3/2 or -1/2

OpenStudy (anonymous):

it would be sqrt2/2 because we know that it has to be positive since it is in the first quadrand and our reference angle is pi/4

OpenStudy (anonymous):

that is what i got if the equation was positive so it is still the same if it is negative?

OpenStudy (anonymous):

not all the time...  the values of cos are positive in the first and 4th quadrant because 7pi/4 takes you to the fourth quadrant and -7pi/4 takes you to the first quadrant the answers will be the same